OpenStudy (anonymous):

Solve: 3 ( (-5) + 7) + ((19-7)/4)

OpenStudy (anonymous):

Solve what is inside the parenthesis first 3(-5+7) + (12/4) = 3(2) + 3 does that make sense?

OpenStudy (anonymous):

the answer is 9.

OpenStudy (anonymous):

Sorry that should be 3(2) + 3(3)

OpenStudy (anonymous):

So the answer is actually 15

OpenStudy (anonymous):

and no its 9. and I'll show you why

OpenStudy (anonymous):

3((-5)+7)+((19-7)/4) 3(2)+(12/4) 3(2)+3 6+3 =9 @KirbyLegs

OpenStudy (anonymous):

Yes but the 3 is multiplied to each term. There is a double parenthesis between the first term, and a closing parenthesis at the end term 3 [ ((-5) + 7) + (19-7)/4]

OpenStudy (anonymous):

Okay, so now I'm a bit confused. Which is it?

OpenStudy (anonymous):

I am not here to fight. I am here to just help. But Mr. @KirbyLegs's percentage is higher.

OpenStudy (anonymous):

Just submit your answer and then come back and tell us what the right answer was Thank you

OpenStudy (anonymous):

I know your not Diamond, and thanks for your help